Solve the differential equation of the following homogeneous equation. Show complete solution.


(3x^2 y - 6x)dx + (x^3 + 2y)dy = 0


1
Expert's answer
2021-09-30T02:07:41-0400

"d(x^3y-3x^2+y^2)=0;\\
x^3\cdot y-3x^2+y^2=C;\\
y^2+x^3\cdot y-3x^2-C=0;\\
y_{1,2}=\frac{-x^3\pm \sqrt{x^6+12x^4+C}}{2},\space C\in R"